find the values of k for which the following equations have two distinct roots.

(a) x^2+8x+3=k

(b) 2x^2-5x =4-k

(c) kx^2-4x+2=0

(d) kx^2+2(k-1)x+k=0

(e) 2x^2=2(x-1)+k

(f) kx^2+(2k-5)x=1-k


1
Expert's answer
2022-01-10T16:50:21-0500

(a)


"x^2+8x+3-k=0"

"D=(8)^2-4(1)(3-k)>0"

"4(13+k)>0"

"k>-13"

"k\in( -13, \infin)"

(b)


"2x^2-5x+k-4 =0"

"D=(-5)^2-4(2)(k-4)>0"

"57-8k>0"

"k<\dfrac{57}{8}"

"k\in(-\infin, \dfrac{57}{8})"

(c)


"kx^2-4x+2=0"

"k\not=0"

"D=(-4)^2-4(k)(2)>0"

"16-8k>0"

"k<2"

"k\in(-\infin, 0)\cup(0, 2)"


(d)


"kx^2+2(k-1)x+k=0"

"k\not=0"

"D=(2(k-1))^2-4(k)(k)>0"

"1-2k>0"

"k<\dfrac{1}{2}"

"k\in(-\infin, 0)\cup(0, \dfrac{1}{2})"

(e)


"2x^2=2(x-1)+k"

"x^2-x+(1-\dfrac{k}{2})=0"

"D=(-1)^2-4(1)(1-\dfrac{k}{2})>0"

"-3+2k>0"

"k>\dfrac{3}{2}"

"k\in (\dfrac{3}{2}, \infin)"

(f)


"kx^2+(2k-5)x+k-1=0"

"k\not=0"


"D=(2k-5)^2-4(k)(k-1)>0"

"4k^2-20k+25-4k^2+4k>0"

"-16k+25>0"

"k<\dfrac{25}{16}"

"k\in(-\infin, 0)\cup(0, \dfrac{25}{16})"